OpenStudy (anonymous):

The probability that Edward purchases a video game from a store is 0.67 (event A), and the probability that Greg purchases a video game from the store is 0.74 (event B). The probability that Edward purchases a video game (given that Greg has purchased a video game) is 0.67. Which statement is true?

OpenStudy (anonymous):

Events A and B are independent because P(A|B) = P(A) Events A and B are dependent because P(A|B) = P(A)

OpenStudy (anonymous):

Events A and B are independent because P(A|B) = P(B) Events A and B are dependent because P(A| B) P(A)